OpenStudy (anonymous):

How do i solve p -4/9p = - 7/9?

OpenStudy (rulnick):

Is it (-4/9) p or -4/(9p)?

OpenStudy (rulnick):

If it's -4/(9p) = -7/9, then cross multiply to get ...

OpenStudy (rulnick):

-4(9) = 9p(-7) -36 = -63p p = 36/63 = ... (reduce)

OpenStudy (rulnick):

p = 4/7

OpenStudy (anonymous):

It is just as I typed it there are no ( )'s

OpenStudy (rulnick):

... so the "p" is in the denominator?

OpenStudy (anonymous):

yes

OpenStudy (rulnick):

Then the solution above is good. You're all set.

OpenStudy (anonymous):

Thanks